A box contains 4 red cubes, 3 blue cubes, 2 green cubes, and 1 yellow cube. You pick two cubes at random, replacing after you pick. Find P(red, then yellow). What's the answer?
Answer by stanbon(75887)

A box contains 4 red cubes, 3 blue cubes, 2 green cubes, and 1 yellow cube. You pick two cubes at random, replacing after you pick. Find P(red, then yellow). What's the answer?
----
Because of replacement the selections are independent.
-----
P(red, then yellow) = (4/10)(1/10) = 4/100
---------------------------
Cheers,
Stan H.
======================